OpenStudy (anonymous):

Two numbers have a sum of 72. What is their product if it is a maximum?

OpenStudy (anonymous):

Well, I think you mean their sum in the second part of the question, and that would just be 72 again.

OpenStudy (dumbcow):

i believe its always a max when numbers are same 36+36 = 72 max product = 36^2

OpenStudy (dumbcow):

heres the reason: x +y = 72 ----> y = 72-x P = xy P = x(72-x) = 72x - x^2 take derivative and set equal to 0 ---> 72 - 2x = 0 x = 36

OpenStudy (anonymous):

half of 72 should do it

OpenStudy (anonymous):

i doubt it is a calculus problem, although i could be wrong

OpenStudy (perl):

36^2

OpenStudy (aum):

Let one number be x. The other number will be (72 - x). Their product y = x * (72 - x) = 72x - x^2 y = -x^2 + 72x This is a parabola that opens downward. Therefore, it attains its maximum at the vertex. The x-coordinate of the vertex of a parabola is given by -b/2a which here is -72/(2*(-1)) = 36. So the product is maximum when x = 36 which makes the other number 72-36=36.